9. Compute the present value of annual perpetual cash flow described as follows. Every odd year (i.e. years 1, 3, 5..) we have a fixed cash flow of $50, and every even year a fixed cash flow of S80 (years 2, 4, 6 ). The annual interest rate is 7%.
Answer #1

present value of annual perpetual cash flow =

we have to find out first present value of effective annual rate of every 2 years

(1+0.07)2 - 1 = 0.1449 = 14.49%

To make sum easy, we find one year future value of 50 = 50*(1+0.07) = 53.5

now 53.5 is receivable at the end of 2nd year

so total amount receivable at the end of 2nd year = 53.5 + 80 = 153.5

to get the answer, we have to find perpetual value of above cashflows

As amount is receivable at the end of 2nd, 4th and so on

we have to use effective annual rate of 2 years = 14.49% (as calculated above)

perpetual value = Amount/Rate

PERPETUAL VALUE = 153.5/0.1449 = 1059.35

ANSWER : $1059.35 (THUMBS UP PLEASE)